# Enhancing Maritime Trajectory Forecasting via H3 Index and Causal Language Modelling (CLM)

### v1

@@ Abstract
The prediction of ship trajectories is a growing field of study in artificial intelligence. Traditional methods rely on the use of LSTM, GRU networks, and even Transformer architectures for the prediction of spatio-temporal series. This study proposes a viable alternative for predicting these trajectories using only GNSS positions. It considers this spatio-temporal problem as a natural language processing problem. The latitude/longitude coordinates of AIS messages are transformed into cell identifiers using the H3 index. Thanks to the pseudo-octal representation, it becomes easier for language models to learn the spatial hierarchy of the H3 index. The method is compared with a classical Kalman filter, widely used in the maritime domain, and introduces the Fréchet distance as the main evaluation metric. We show that it is possible to predict ship trajectories quite precisely up to 8 hours with 30 minutes of context. We demonstrate that this alternative works well enough to predict trajectories worldwide.
@@ 1 Introduction
In light of these considerations, we set out to explore an alternative solution: the discretisation method utilising Uber’s Hexagonal Hierarchical Spatial Index (H3) index [17]. H3 is a geospatial indexing system that partitions the world into hexagonal cells. H3 cells are written to 64 bits and can be represented in hexadecimal format, providing a creative solution to the challenges associated with the continuous representation of GNSS coordinates. Our central aim with this discretisation approach is to reconfigure a ship’s trajectory from Automatic Identification System (AIS)  [^1] The Automatic Identification System is an automatic tracking system used to monitor and locate ships in real time. It uses Very High Frequency (VHF) transceivers to transmit data such as the vessel’s identifier, position, speed and direction. messages [18] into a structured narrative, akin to the composition of words in a sentence. By doing so, we effectively counteract the potential detrimental impacts of minor coordinate inaccuracies, and we expect to benefit from the ability of Transformer architectures to understand the relationships between tokens.
Given the limited usefulness of the initial 12 bits of the H3 cells for improving our understanding of space, we propose an innovative representation. Our aim here is to optimise the cell representation, driven by the need to alleviate computational overhead and establish a more streamlined tokeniser vocabulary, consisting of just 270 tokens. This tailored approach not only enhances the efficiency of processing ship trajectories but also seamlessly aligns them with Transformer-based models equipped with spatial awareness, as envisioned in this research.
In the forthcoming sections, we delve into the methodologies adopted, with a primary emphasis on the utilisation of the H3 index for spatial representation. We elaborate on the rationale behind this choice and its potential to enhance ship trajectory prediction models. Furthermore, we present experimental results and validate our approach against existing methods, demonstrating the promise of integrating Transformer architectures with spatial awareness within the maritime domain.
@@ 2 Background
GNSS play a pivotal role in modern artificial intelligence, particularly in applications related to Geospatial Intelligence (GEOINT). GEOINT, as defined by the USA, refers to the analysis and exploitation of geospatial information to gain insights and support joint operations [19, 20]. GNSS, which includes systems like GPS, GLONASS, and Galileo provides essential geospatial data that underpins these intelligence operations.
@@ 4.1 H3 Index
The H3 index offers a host of advantages ideally suited for our objectives. H3 represents geographical locations as hexagons as shown in figure 2, ensuring that any two neighbouring hexagons are always at an absolute distance of 1 unlike the S2 representation system [35]. The geometric nature of hexagons makes them less deformable at the poles. Furthermore, H3 provides multiple resolution levels, making it adaptable to various spatial scales. The 64-bit cell identifiers in H3 proved highly efficient for handling extensive geographical data, aligning seamlessly with the demands of big data applications. Unlike alternative methods [36], such as Geohash and Hexbin, H3 demonstrated reduced visual distortions when subjected to different map projections, ensuring the preservation of spatial fidelity.
@@
Tokenisation played a pivotal role in managing our spatial representation. We considered each two-character basic cell as a single token and treated each resolution level as a distinct token. Since we chose resolution 10, all informative tokens after this resolution are non-relevant and were thus set to 7 (empty octal value, resolutions 11 to 15). We could have removed them or tokenised as a single token to reduce computational complexity at the cost of a more complex return to coordinates in longitude and latitude. However, we chose to maintain them in our methodology, and we discuss its implications for future refinements.
@@
4.2 Model Architecture: Mixtral8x7B
Our research leverages the powerful architecture provided by the Mixtral8x7B model [37]. This model is based upon the Mistral 7B architecture [38] and incorporates a Mixture of Experts approach [39, 40], significantly enhancing its capacity to handle complex tasks. The inclusion of a sliding window mechanism further sets it apart, allowing for an attention span of 2560 tokens—a feature particularly beneficial when dealing with lengthy trajectories. In our specific case, this translates to an attention span of over 2 hours, providing the ability to capture intricate patterns in maritime movements. Additionally, the utilisation of Grouped Query Attention not only facilitates faster inference but also reduces cache size requirements, ensuring efficient model operation.
@@ 4.2 Model Architecture: Mixtral8x7B
4.3 Hardware, dataset and training process
@@
The model was trained on a machine equipped with an Intel Core i9-13900K processor, an NVIDIA GeForce RTX 4090 graphics card with 24GB of RAM on a Manjaro 23.1.3 operating system. The model was implemented using the HuggingFace and the PyTorch frameworks. The training of the model took 175,000 steps using an Adam optimiser with an initial learning rate of 5 times $10^{-4}$ and a gradient accumulation of 8 (see B for the learning curves).
The dataset used to train the model comes from a capture we made ourselves and covers a period of one week in August 2023. The working area forms a rectangle with latitude/longitude coordinates $(-43.33,-88.59;\quad\allowbreak 80.25,168.34)$, but the majority of the trajectories are located in the North Sea, English Channel, Atlantic Ocean and Mediterranean Sea.
Tokenised trajectories exceeding the size of the model context are split into chunks of size 2560. We choose to keep the overlapping tokens and to set a padding with an attention mask to zero. It consists of 1,731,686 unique trajectories within a set of around 32 million H3 cells. After tokenisation, it consists of 4 billion tokens and, to date, we believe it is the largest set of trajectory data cleaned and prepared for trajectory forecasting. The entire captured dataset has been used for training and validation (respectively 85% and 15% of the dataset). The testing phase was carried out using another dataset that we captured on the fly during 3 days in February 2024 and represents approximately 15,000 cleaned trajectories.
@@ 4.4 Evaluation metric
To analyse the results, we propose utilising the Fréchet distance [41] given by the equation LABEL:eq:frechet. The rationale behind this choice lies in its ability to capture the similarity between two trajectories while considering both spatial proximity and sequential order of points. The Fréchet distance, by comparing the geometric similarity between the actual and predicted trajectories, serves as a stringent measure of predictive fidelity. We implemented the algorithm using the Haversine function, a method commonly employed for calculating distances between points on the Earth’s surface due to its accuracy in spherical geometry.
@@
4.5 Kalman filter
To the best of our knowledge, there is currently no method for predicting the trajectory of ships at any point on Earth and over a long distance without being adjusted to the actual observations that will be made during data capture. We have chosen to compare our method with Kalman filters [42]. Our model uses only GPS positions, specifically longitude (lon) and latitude (lat), to predict trajectories, and a prediction is represented by a matrix $G\in\mathbb{R}^{\tau\times 2}$ with $\tau$ the length of the prediction. In order to more closely align with our model, which utilises only the positions for predictions, we chose to include only four values in the state of the Kalman filter: the longitude (lon), the latitude (lat), and their velocities ($v_{lon}$ and $v_{lat}$). Traditionally, Extended Kalman filters can include more complex data like sea current patterns. However, to reduce dependency to external data, we chose to reduce the number of parameters for our model and thus for the Kalman filter. The Kalman filter predicts a single state and must adapt to real observations in its update phase. However, real observations do not yet exist in our maritime case. We need to modify the Kalman filter to make it auto-regressive with a number of steps $\tau$, considering that a predicted state becomes the truth. We have chosen to restrict ourselves to the prediction phase, as expressed in the following equation LABEL:eq:kalman.

### v2

@@ Abstract
The prediction of ship trajectories is a growing field of study in artificial intelligence. Traditional methods rely on the use of LSTM, GRU networks, and even Transformer architectures for the prediction of spatio-temporal series. This study proposes a viable alternative for predicting these trajectories using only GNSS positions. It considers this spatio-temporal problem as a natural language processing problem. The latitude/longitude coordinates of AIS messages are transformed into cell identifiers using the H3 index. Thanks to the pseudo-octal representation, it becomes easier for language models to learn the spatial hierarchy of the H3 index. The method is compared with a classical Kalman filter, widely used in the maritime domain, and introduces the Fréchet distance as the main evaluation metric. We show that it is possible to predict ship trajectories quite precisely up to 8 hours ahead with 30 minutes of context, using solely GNSS positions, without relying on any additional information such as speed, course, or external conditions — unlike many traditional methods. We demonstrate that this alternative works well enough to predict trajectories worldwide.
@@ 1 Introduction
In light of these considerations, we set out to explore an alternative solution: the discretisation method utilising Uber’s Hexagonal Hierarchical Spatial Index (H3) [17]. H3 is a geospatial indexing system that partitions the world into hexagonal cells. H3 cells are written to 64 bits and can be represented in hexadecimal format, providing a creative solution to the challenges associated with the continuous representation of GNSS coordinates. While the concept of partitioning trajectories into hexagonal cells has already been explored by Spadon et al. [18], their approach does not strictly adhere to Uber’s original H3 method. Instead, they propose a similar discretisation technique, which inspired us to refine the use of H3 for our purposes. Our central aim with this discretisation approach is to reconfigure a ship’s trajectory from Automatic Identification System (AIS)  [^1] The Automatic Identification System is an automatic tracking system used to monitor and locate ships in real time. It uses Very High Frequency (VHF) transceivers to transmit data such as the vessel’s identifier, position, speed, and direction. messages [19] into a structured narrative, akin to the composition of words in a sentence. By doing so, we effectively counteract the potential detrimental impacts of minor coordinate inaccuracies, and we expect to benefit from the ability of Transformer architectures to understand the relationships between tokens.
Given the limited usefulness of the initial 12 bits of the H3 cells for improving our understanding of space, we propose an innovative pseudo-octal format. This representation is a cornerstone of our study. By tailoring this and integrating it with a customised tokenisation scheme, we have constructed a tokeniser vocabulary of just 270 tokens. This streamlined encoding not only optimises the computational load but also facilitates the Transformer models’ spatial comprehension. Together, this pseudo-octal representation and tokenisation allow the models to grasp the structure of ship trajectories, offering a significant advantage in spatial pattern recognition within our Transformer-based architecture.
To analyse the results, we propose utilising the Fréchet distance [20]. The rationale behind this choice lies in its ability to capture the similarity between two trajectories while considering both spatial proximity and sequential order of points. The Fréchet distance, by comparing the geometric similarity between the actual and predicted trajectories, serves as a stringent measure of predictive fidelity.
In the forthcoming sections, we delve into the methodologies adopted, with a primary emphasis on the utilisation of the H3 index for spatial representation. We elaborate on the rationale behind selecting H3 at a zoom level of 10, which directly influences the granularity of the discretised ship trajectories. This choice plays a critical role in the quality of predictions, as the level of precision in the H3 index determines the amount of zigzagging introduced into a trajectory after discretisation. A finer resolution could capture more detailed movements, but may also increase noise, whereas a coarser resolution might miss important spatial nuances.
Furthermore, we present experimental results and validate our approach against existing methods, demonstrating the promise of integrating Transformer architectures with spatial awareness within the maritime domain.
@@ 2 Background
GNSS play a pivotal role in modern artificial intelligence, particularly in applications related to Geospatial Intelligence (GEOINT). GEOINT, as defined by the USA, refers to the analysis and exploitation of geospatial information to gain insights and support joint operations [21, 22]. GNSS, which includes systems like GPS, GLONASS, and Galileo, provides essential geospatial data that underpins these intelligence operations.
@@ 3 Related Work
A recent study by Spadon et al. [18] introduces a novel approach to enhance vessel trajectory forecasting through the integration of probabilistic modeling and advanced neural network techniques. The authors made significant contributions to maritime safety through vessel trajectory forecasting. One key innovation is the use of a probabilistic model that leverages historical AIS data to create Route Probability and Destination Probability Matrices. These matrices generate probable routes and destinations, adding valuable features to the vessel trajectory data. This approach significantly enhances the model’s decision-making capacity, particularly for complex or curved paths, by reducing mean and median errors. The integration of these probabilistic features allows the model to distill generalized patterns and rules from the data, making it more robust and adaptable to various conditions.
Another notable aspect is the positional-aware attention mechanism, which prioritizes more recent time steps of the input sequence. This methodological focus ensures that recent data points have a greater impact on the learning process, enhancing the precision of the trajectory reconstruction task. By giving higher importance to the latest time steps, the model can capture nuanced temporal patterns that improve its trajectory-fitting capabilities. This approach allows the model to navigate the complexity of maritime navigation more effectively, making it well-suited for real-time applications processing streaming AIS data.
The authors’ methodology emphasises the significance of recent data points in the learning process. However, similar to the TrAISformer model [35], the dataset used in this study is limited to a specific geographical region. Although the objective is not to generalise across all global trajectories, the application of a probabilistic model to enhance learning indicates that the transferability of this model to other regions may require recalculating the probability matrix.
@@ 4.1 H3 Index
The discretisation of a near-continuous space such as the GNSS coordinate system [^2] Although the GNSS coordinate system is technically discrete, with a smallest measurable unit of 1 mm, we refer to it as ”near-continuous” because this level of precision is negligible compared to the scale of decimal degrees. For instance, the first digit in decimal degrees corresponds to a precision of approximately 111 km [37]., through the use of the H3 index is a rational choice for mitigating inaccuracies in geographical coordinates. By representing spatial data as discrete hexagons, the H3 index ensures that small deviations in coordinates do not translate into significant errors. Each hexagon serves as a well-defined spatial unit, reducing the sensitivity of measurements to minor fluctuations in GNSS data. This contrasts with continuous coordinate systems, where slight variations can lead to substantial discrepancies. Furthermore, the multiple resolution levels of H3 allow for adaptability across different spatial scales, ensuring both accuracy and computational efficiency. This approach not only alleviates the impact of measurement uncertainties but also enhances the robustness of spatial data representation, making it a practical solution for maritime trajectory prediction.
In addition to its robustness, the H3 index offers several advantages that align well with our objectives. H3 represents geographical locations as hexagons, ensuring that any two neighbouring hexagons are always at an absolute distance of 1, unlike the S2 representation system [38]. The geometric nature of hexagons makes them less deformable at the poles, and H3 provides multiple resolution levels, making it adaptable to various spatial scales. The 64-bit cell identifiers in H3 are highly efficient for handling extensive geographical data, aligning seamlessly with the demands of big data applications. Unlike alternative methods [39], such as Geohash and Hexbin, H3 demonstrates reduced visual distortions when subjected to different map projections, ensuring the preservation of spatial fidelity.
@@
The proposed algorithms for converting H3 cell identifiers to and from a pseudo-octal representation are crucial for optimising the spatial encoding process and addressing limitations in traditional representations. The default hexadecimal representation of the H3 index is inadequate for our use case, as it merely translates the cell identifier without encoding spatial characteristics. This makes it cumbersome for our tokenisation process, as it does not map directly to the hierarchical structure of H3 cells. Likewise, using the 64-bit binary representation would necessitate extremely long sequences for each position, limiting the efficiency of trajectory predictions. In contrast, the pseudo-octal representation leverages the hierarchical nature of the H3 index, where each subsequent resolution is formed using a 7-cell aperture. By discarding the initial 12 bits and encoding each resolution level in octal form (except the base cell, which remains in hexadecimal), we achieve a compact and efficient representation aligned with the spatial hierarchy of the H3 index. This encoding scheme is particularly valuable for our application in predicting maritime trajectories, where optimising sequence length enhances the model’s ability to predict trajectories accurately and efficiently.
@@
Tokenisation plays a pivotal role in managing this spatial representation within our text-based model. Each two-character base cell is treated as a single token, with every subsequent resolution level encoded as a distinct token. Since we use resolution 10, all tokens beyond this resolution (levels 11 to 15) are non-informative and set to 7 (an empty octal value). While we could have removed these or tokenised them as a single token to reduce computational complexity, we opted to retain them in our methodology for now. This leaves open the possibility of revisiting the tokenisation in the future, should we choose to adjust the resolutions without modifying the algorithms.
@@
4.2 Model Architecture
In the pursuit of predicting ship trajectories, the selection of an appropriate model architecture is crucial. This section delves into the rationale behind choosing a causal language model and explores the intricacies of the Mixtral8x7B framework. Causal models, with their auto-regressive nature, are particularly suited for tasks that demand sequential prediction, making them ideal for capturing the temporal and causal structure of the trajectories. This section will detail the specific configuration of our architecture, highlighting its capacity and flexibility in handling the complexities of maritime trajectory prediction.
4.2.1 Rationale for Selecting a Causal Model
The choice of a causal language model, also referred to as an auto-regressive model, is grounded in its inherent ability to predict future tokens based solely on preceding ones [40, 9, 8]. This auto-regressive framework aligns naturally with many linguistic tasks, as language generation is inherently sequential. In CLM, each word or token is predicted based on all the prior tokens, capturing the temporal and causal structure of language. This makes CLMs particularly effective for tasks where the generation of each token depends on the context provided by the previous ones (unidirectional, left-to-right in our case).
One of the primary strengths of causal models is their ability to handle long-range dependencies within a sequence. Given that language often exhibits dependencies between distant elements, CLMs are well-suited for capturing these relationships over extended sequences. As a result, they can generate text that remains coherent and contextually appropriate over longer spans, which is critical for many downstream applications like our case.
In the context of predicting ship trajectories, the use of a causal language model is particularly justified by the sequential nature of vessel movement, where each GNSS coordinate is dependent on prior positions. By transforming each coordinate into a pseudo-octal representation, we effectively discretise the spatial domain, capturing the hierarchical and spatial characteristics of each location. The auto-regressive nature of the causal language model ensures that each subsequent position is predicted based on the previous sequence of H3 cells, maintaining consistency in the spatial path. This approach is particularly robust in scenarios where long-range dependencies must be captured—such as forecasting future movement over extended distances—while the structured encoding facilitates the model’s understanding of spatial continuity.
4.2.2 Mixtral8x7B
Our research leverages the powerful architecture provided by the Mixtral8x7B model [41]. This model is based upon the Mistral 7B architecture [42] and incorporates a Mixture of Experts approach [43, 44], significantly enhancing its capacity to handle complex tasks. Each expert can focus on different part of the pseudo-octal representation, which improves predictive accuracy by enabling more granular analysis.
The inclusion of a sliding window mechanism further sets it apart, allowing for an attention span of 2560 tokens—a feature particularly beneficial when dealing with lengthy trajectories. In our specific case, this translates to an attention span of over 2 hours, providing the ability to capture intricate patterns in maritime movements. Additionally, the utilisation of Grouped Query Attention not only facilitates faster inference but also reduces cache size requirements, ensuring efficient model operation.
Additionally, Mixtral’s integration with the HuggingFace Transformers library [45] made it an appealing option for development. This compatibility simplifies the model’s implementation and ensures that our solution is maintainable and extensible [^3] At the time of the revision of this article, more recent models have been released, and we have not reconducted the experiments with these newer models..
@@ 4.2.2 Mixtral8x7B
4.3 Dataset, hardware and training process
@@
The model was trained on a machine equipped with an Intel Core i9-13900K processor, an NVIDIA GeForce RTX 4090 graphics card with 24GB of RAM on a Manjaro 23.1.3 operating system. The model was implemented using the HuggingFace and the PyTorch frameworks.
4.3.1 Dataset and Preprocessing
The AIS data utilised in this study were collected over a period of one week in August 2023, within a defined geographical region spanning latitude/longitude coordinates from $(-84.88,-180.00;\quad\allowbreak 90.00,179.65)$ as shown in figure 3. While the collection area covers a wide range, the majority of the recorded vessel trajectories are concentrated in the North Sea, the English Channel, the Atlantic Ocean (along the French, Portuguese, and Spanish coastlines), and the Mediterranean Sea.
As highlighted in the introduction, AIS data are derived from GNSS coordinates, which may suffer from inaccuracies. Common issues encountered included erratic jumps in position, sudden “teleportations” of points, and multiple vessels sharing the same Maritime Mobile Service Identity (MMSI), despite the AIS standard requiring MMSIs to be unique. While the protocol itself stipulates unique MMSI assignments, this is not strictly enforced, either technically or by software, leading to potential confusion in trajectory analysis.
To address these challenges, we employed several data cleaning and clustering techniques. The first two issues—erratic points and teleportation—were resolved using the Density-Based Spatial Clustering of Applications with Noise (DBSCAN) algorithm, as described by Ester [46], with specific optimisations implemented by Wang [47]. We set the DBSCAN epsilon parameter to 10, which filtered out points that did not belong to any meaningful trajectory, as isolated points were considered noise.
For the issue of duplicate MMSIs, a more intricate process was required. The data were first grouped by MMSI, and for each group, we implemented a systematic cleaning procedure. We began by removing duplicate positions and timestamps, ensuring only unique points remained. Additionally, any point where the Speed Over Ground (SOG) was recorded as zero was discarded, as such records are typically associated with stationary vessels.
Following this, the data were sorted chronologically, and a secondary DBSCAN with an epsilon value of 0.1 was applied. This allowed for clustering of spatially related points, with clusters representing coherent vessel trajectories. Any points labelled as noise by the algorithm were excluded. Subsequently, we computed the centroid of each cluster, as well as the average distance from the cluster’s points to this centroid. Clusters with an average point-to-centroid distance of less than 15 metres were considered artefacts of GNSS imprecision and were removed, as they likely represented stationary vessels (for example, a docked ship).
At this stage, the remaining clusters represented valid trajectories. To detect instances where multiple vessels might share the same MMSI, we examined the temporal gaps between consecutive points, assigning a new group identifier whenever a time difference exceeding five hours was detected. A unique hash was generated for each point based on the cluster and group number, facilitating the identification of distinct vessels using the same MMSI.
We further refined the trajectories by calculating the Haversine distance between consecutive points and estimating the vessel’s actual speed, independent of the recorded SOG. Any points indicating a speed exceeding 100 km/h were flagged as "teleportations" [^4] Given that the theoretical maximum speed is 188.904 km/h (meaning that the vessel or helicopter is travelling at or above this speed)., and the entire trajectory containing such points was discarded, based on the actual hash. Consecutive clusters were then iteratively merged by comparing the final point of one cluster with the initial point of the next. If the transition was feasible based on the average speed of the last 10 points (within a ±$\pm 5\%$ tolerance), the two clusters were combined into a single trajectory, with cluster B inheriting the hash and group number of cluster A.
Finally, a consolidated hash was generated for each trajectory, incorporating the MMSI, trajectory number, and individual cluster hash. Any resulting trajectories with fewer than 10 points were excluded from further analysis to maintain data integrity and avoid bias from insufficiently sampled paths. To standardise the temporal resolution, we then applied a downsampling interval of one minute, ensuring consistency across all trajectories.
In addition to the trajectory refinement, the next step involved transforming the AIS points into spatially encoded identifiers using the H3 indexing system. As described earlier in the article, this transformation followed the procedures outlined in algorithms 1 and 2, effectively converting the vessel locations into hexagonal cells for more efficient spatial representation.
Once transformed, the tokenisation process was applied to prepare the data for trajectory forecasting. Each unique trajectory was tokenised, with the resulting sequence split into chunks of size 2560 whenever a trajectory exceeded the context size of the model. Overlapping tokens were preserved, and padding was introduced with an attention mask set to zero. This approach allowed us to handle large datasets without compromising model accuracy. The final dataset comprises 1,731,686 unique trajectories spread across approximately 32 million H3 cells. After tokenisation, it consists of 4 billion tokens.
This multi-step approach, incorporating spatial clustering, temporal segmentation, H3 transformation, tokenisation, and trajectory filtering, successfully mitigated issues of duplicate MMSIs and erroneous data points, ensuring a high-quality dataset for subsequent analysis and forecasting.
Figure 3: Comparison of maritime trajectories before and after dataset cleaning. (a) Before cleaning, the trajectories exhibit numerous errors, with irregular jumps crossing over land and a high concentration of anomalies covering Europe. These errors indicate a poor capture of the actual ship movements. (b) After applying our cleaning procedure, the trajectories are corrected and closely follow the maritime contours of the continents, removing artefacts and errors that cross landmasses. This process ensures a more accurate representation of the actual maritime routes.
4.3.2 Training and Validation
The training of the model took 175,000 steps using an Adam optimiser with an initial learning rate of 5 times $10^{-4}$ and a gradient accumulation of 8 (see figure 10 in appendix B for the learning curves).
The entire captured dataset has been used for training and validation (respectively 85% and 15% of the dataset). The testing phase was carried out using another dataset that we captured on the fly during 3 days in February 2024 and represents approximately 15,000 cleaned trajectories.
@@ 4.4 Evaluation metric
To further enhance the evaluation of our trajectory prediction models, we employ the Fréchet distance as a key metric given by the equation LABEL:eq:frechet. This choice is crucial in addressing the limitations of traditional error metrics such as Mean Absolute Error (MAE) or Mean Squared Error (MSE) [35, 18], which compare points in isolation and assume a strict correspondence between them, potentially overlooking the overall shape and progression of the trajectory. The Fréchet distance, by contrast, evaluates the entire trajectory, accounting for both the geometry and the order of points, making it better suited for spatial and temporal analysis. This allows us to capture the continuity and structure of maritime movements more effectively, providing a more comprehensive assessment of model performance as shown in figure 4. We implemented the algorithm using the Haversine function, a method commonly employed for calculating distances between points on the Earth’s surface due to its accuracy in spherical geometry.
Figure 4: Comparison of the Fréchet distance versus Mean Absolute Error (MAE) and Mean Squared Error (MSE). Over a distance of 596 km, the predicted trajectory deviates significantly from the ground truth, making such a prediction unacceptable. The MAE gives the lowest error, but this value is misleading as it does not capture the true nature of the deviation, being too low due to its point-by-point comparison. In contrast, the Fréchet distance accounts for the overall shape of the trajectories, providing a higher and more realistic measure of error.
@@
4.5 Qualitative and Quantitative Evaluation of Trajectory Predictions
To the best of our knowledge, there is currently no method capable of predicting ship trajectories globally and over long distances without being tailored to specific geographic regions. Most existing approaches are limited to particular areas, restricting their applicability on a broader scale. In contrast, we sought to develop a solution that is not constrained by these regional limitations.
For the qualitative evaluation on our test dataset, we have chosen to compare our approach with Kalman filters [48]. This widely used method is known for its adaptability in trajectory prediction, yet it still requires regional adjustments, making it an apt baseline to highlight the advantages of our model. In our implementation, only GPS positions—longitude (lon) and latitude (lat)—are used to predict trajectories. Each prediction is represented by a matrix $G\in\mathbb{R}^{\tau\times 2}$, where $\tau$ is the length of the prediction. To align with our approach, which relies solely on positional data, we restricted the state in the Kalman filter to include only four values: longitude (lon), latitude (lat), and their respective velocities ($v_{lon}$ and $v_{lat}$). Although Extended Kalman filters can incorporate more complex data such as sea current patterns, we opted to reduce dependency on external data by limiting the parameters of both our model and the Kalman filter. Since the Kalman filter predicts a single state and relies on real observations in its update phase, and real observations are not yet available in our maritime case, we modified the Kalman filter to operate in an auto-regressive mode over $\tau$ steps. In this setup, each predicted state becomes the subsequent input in the prediction cycle, as expressed in equation LABEL:eq:kalman.